OpenStudy (anonymous):

write the equation of the line passing through (-3,2) and is parallel to and perpendicular to the line y=6x-3

OpenStudy (anonymous):

is it parallel or perpendicular

OpenStudy (anonymous):

????????????????????????????

OpenStudy (anonymous):

both write one for being perpendicular and one of being parallel

OpenStudy (anonymous):

PLEASE AND THANKYOU :)

OpenStudy (anonymous):

for perpendicular: the slope of the line is -1/6 y = -(1/6)x +c 2 = -1/6(-3) +c c= 3/2 =>line: y=-(1/6)x+3/2

OpenStudy (anonymous):

for parallel: the slope of the line is 6 =>y=6x+c =>2=-18+c =>c=20 =>line: y=6x+20

OpenStudy (anonymous):

It's incorrect how did you calculated the c in perpendicular line.

OpenStudy (anonymous):

thats what i was giong to ask

OpenStudy (anonymous):

when it's perpendicular the equation must be y = -1/6x - 8/3

OpenStudy (anonymous):

i dont know where c even came from haha and could you explain how you got the answer and okay thankyou

OpenStudy (anonymous):

where did the eight and three come from ugh!!

OpenStudy (anonymous):

if the line is perpendicular the slope will be - 1/6, and so substitute the value of slope and point in point slope equation, manipulate it and get it in slope intercept.

OpenStudy (anonymous):

thats what i need help with thats what i dont know how to do haha

OpenStudy (anonymous):

k wait i will draw it.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

I am extremly sorry the equation of perpendicular line will be y = -1/6x - 3/2

OpenStudy (anonymous):

dont worry about it that was great help i know who to go to if i need help!! :) :)

OpenStudy (anonymous):

for parallel line : y - 6x + 20